# Multi-Group Proportional Representation

### v1

@@ Appendix C Description of Fair Retrieval Algorithms and Retrieval Statistics and Other Related Works
In this section we include additional experimental results. First, we include full tables of intersectional retrieval performance on UTKFace for our method, $k$-NN, MMR, ClipClip, PBM, and DebaisCLIP. We average over 10 queries as while retrieving 50 images. In Table 3, we compute MPR with a linear regression oracle and optimize our method with the same linear regression oracle. We are able to perfectly balance across intersectional groups while other methods often fail to account for the ‘Others’ race category and the ‘Female’ gender category. Some methods such as PBM which optimize for equal representation are able to achieve balanced representation across the gender axis but fail to consider diversity in race. In Table 4, we see a similar story for MPR measured and optimized with a decision tree.
@@
Next, we include additional similarity-MPR tradeoff curves when measuring and optimizing MPR with a decision tree (Figure 10) and with an MLP with 1 hidden layer of dimension 64 (Figure 11). We observe a similar Pareto-domination to that observed in the main paper, as our method is able to lower MPR the most while preserving similarity even with more nonlinear oracle estimators of MPR. Figure 11, we plot on a log-scale for MPR as the MLP is able to find such violating functions for baseline methods that their “Fraction of top-k MPR" can reach values of 35. In the MLP setting, while PBM preserves high similarity on a whole, it is still unable to decrease MPR as low as MAPRis able to at $\approx$ 95% of top-k similarity.
In Figure 12 we include nine additional precision-representation curves. Similar to the result for the query “chief executive officer" in the main paper, MAPR maintains similar precision-representation performance to baseline methods.
@@
Finally, we include example retrievals for ‘architect’ and ‘teacher’ queries over UTKFace with each of our baselines. Similar to the tables, we select each algorithm’s solution with the lowest MPR after doing a hyperparameter sweep for each respective algorithm and plot the corresponding images below.
@@
Table 3: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [39] for MAPR, $k$-NN, MMR [27], ClipClip PBM, and DebiasCLIP with MPR optimized and measured with a linear regression oracle. Entries indicate average percentage representation in retrieved items. MAPR is able to balance representation across classes, whereas other baselines miss intersectional groups (highlighted in red).
@@
MAPR  (Ours) Male 7.8 $\pm$ 3.74 10.2 $\pm$ 2.90 13.0 $\pm$ 2.56 11.8 $\pm$ 2.44 7.2 $\pm$ 3.70 50 $\pm$ 0
@@
Table 4: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [39] for MAPR, $k$-NN, MMR [27], PBM, and DebiasCLIP with MPR optimized and measured with a decision tree oracle. Entries indicate average percentage representation in retrieved items. MAPR is able to balance representation across classes, whereas other baselines miss intersectional groups (highlighted in red).
@@
Figure 10: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR for MPR measured by function class $\mathcal{C}$ given by the set of decision trees. Averaged over 10 queries for $k=50$ images retrieved. From Left-to-Right: CelebA, UTKFaces, Occupations. Values are normalized so $k$-NN achieves point (1,1) in each case. MAPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 11: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR for MPR measured by function class $\mathcal{C}$ given by the set of 2-layer MLPs with hidden dimension 64. Averaged over 10 queries for $k=50$ images retrieved. From Left-to-Right: CelebA, UTKFaces, Occupations. Values are normalized so $k$-NN achieves point (1,1) in each case. MAPR Pareto-dominates baselines and significantly closes the MPR gap.
@@
Figure 13: Retrievals for MAPR with query “A photo of a teacher"
@@
Figure 16: Retrievals for ClipClip with query “A photo of a teacher"
@@
Figure 19: Retrievals for MAPR  with query “A photo of an architect"
@@
Figure 22: Retrievals for ClipClip with query “A photo of an architect"
@@
Appendix E Experiment Notes
E.1 GPT Generated Queries.
@@ Appendix E Experiment Notes
E.2 Experiment Setting and Details.
These experiments were conducted mainly on cpu after initially using a single A100 GPU to compute CLIP and DebiasCLIP embeddings and run FAISS to compute the top 10k candidates for each. Then, all embeddings are saved so that retrieval across all methods can be run on cpu. When training linear probes, we conduct cross-validated hyperparameter search from $0.01$ to $100$ over $\ell_{2}$ regularized logistic regression modesl from scikit-learn [68] to find the optimal probe.
To run MAPR, we use Gurobi [90] with an Academic License to solve Eqn. (10) each iteration of our cutting plane method, and use CVXPY [91, 92] to solve our closed-form, quadratic program.
@@
On average, the runtime for MAPR is approximately 1-5 minutes with 50 queries and 10000 elements in each of our retrieval and curation sets. However, with more features (such as using CLIP embeddings instead of group attributes) this runtime can increase. We note however, that our most competitive baseline, MMR, is a greedy algorithm that requires 1-2 hours per query to compute. In live image retrieval settings, this runtime is too slow to be of any practical use.

### v2

@@ Appendix G Additional Results
Second, we include full tables of intersectional retrieval performance on UTKFace for our method, Top-$k$, MMR, CLIP-Clip, PBM, and DebiasCLIP. We report averaged results for each of the 10 queries while retrieving 50 images. In Table 5, we compute MPR with a linear regression oracle and optimize our method with the same linear regression oracle. We are able to perfectly balance across intersectional groups while other methods often fail to account for the ‘Others’ race category and the ‘Female’ gender category. Some methods, such as PBM, which optimize for equal representation, are able to achieve balanced representation across the gender axis but fail to consider diversity in race. In Table 6, we see a similar story for MPR measured and optimized with a decision tree.
@@
Next, we include additional similarity-MPR tradeoff curves for the other 9 queries not reported in the main paper (Figs. 8,9,10), as well as for all 10 queries on two additional models: a decision tree (Figs. 11, 12, 13) and with an MLP with 1 hidden layer of dimension 64 (Figs. 14, 15, 16). We observe a similar Pareto-domination to that observed in the main paper, as our method is able to lower MPR the most while preserving similarity even with more nonlinear oracle estimators of MPR. In the MLP setting, while PBM preserves high similarity on the whole, it is still unable to decrease MPR as low as MOPRis able to at $\approx$ 95% of top-k similarity.
We also include results for an additional 45 queries for the Occupations dataset (with FairFace as the curation dataset), as this dataset comes with 45 ground-truth occupations labels. These results again highlight the Pareto-dominance of MOPR, as shown in Figures 17, 18, and 19.
In Figure 20, we explore the performance of MOPR in terms of precision, a common retrieval metric. We can do this using the Occupations dataset, as it has ground truth occupations labels with which we can benchmark retrieval precision. We see that MOPR performs competitively in terms of precision while also traversing more of the precision-MPR Pareto frontier. In Figure 21, we include nine additional precision-representation curves. Similar to the result for the query “chief executive officer", MOPR maintains similar precision-representation performance to baseline methods.
@@
Finally, we include example sets of retrievals for ‘architect’ and ‘teacher’ queries over UTKFace with each of our baselines. Similar to the tables, we select each algorithm’s solution with the lowest MPR after doing a hyperparameter sweep for each respective algorithm and plot the corresponding images below. The rest of the Appendix is solely tables and figures.
k=10 k=50 k=150
MOPR(LP, 10 iterations) 1.434 $\pm$ 0.304 1.6 $\pm$ 0.54 1.371 $\pm$ 0.361
MOPR(LP, 50 iterations) 5.14 $\pm$ 0.352 4.199 $\pm$ 1.74 4.342 $\pm$ 1.803
MOPR(QP, linear regression) 0.528 $\pm$ 0.061 0.533 $\pm$ 0.066 0.584 $\pm$ 0.089
MMR 151.961 $\pm$ 1.126 267.412 $\pm$ 3.863 556.184 $\pm$ 6.324
PBM 0.275 $\pm$ 0.162 0.241 $\pm$ 0.013 0.276 $\pm$ 0.151
CLIP-Clip 21.303 $\pm$ 0.361 21.907 $\pm$ 0.525 21.34 $\pm$ 0.353
DebiasCLIP 2e-4 $\pm$ 1e-6 2e-4 $\pm$ 1e-6 2e-4 $\pm$ 1e-6
Table 3: Average running time (in seconds) over ten queries for k=10, 50, and 150 items for the different methods. MOPR is able to achieve 100x speedup over the SOTA retrieval algorithm, MMR, and maintains competitive performance with vanilla KNN-based methods such as CLIP-Clip and DebiasCLIP. For more discussion, see the general comment in the rebuttal.
Normalized Mean Cosine Similarity
k-NN 1.$0$$\pm$ 0.0000
MOPR 0.9049 $\pm$ 0.0692
MMR 0.8047 $\pm$ 0.1477
PBM 0.9416 $\pm$ 0.0802
CLIP-Clip 0.7758 $\pm$ 0.1433
DebiasCLIP 0.8497 $\pm$ 0.0985
Table 4: Normalized mean cosine similarity between the query embedding and retrieved image embeddings for retrieved images from Tables 1 and 5. The normalization is performed by dividing the mean similarity by the mean similarity given by the Top-$k$ nearest neighbors of a query.
@@
Table 5: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [40] for MOPR, $k$-NN, MMR [27], CLIP-Clip, PBM, and DebiasCLIP with MPR optimized and measured with a linear regression oracle. Entries indicate the average percentage representation in retrieved items. MOPR is able to balance representation across classes, whereas other baselines miss intersectional groups (highlighted in red).
@@
MOPR  (Ours) Male 7.8 $\pm$ 3.74 10.2 $\pm$ 2.90 13.0 $\pm$ 2.56 11.8 $\pm$ 2.44 7.2 $\pm$ 3.70 50 $\pm$ 0
@@
Table 6: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [40] for MOPR, $k$-NN, MMR [27], PBM, and DebiasCLIP with MPR optimized and measured with a decision tree oracle. Entries indicate the average percentage representation in retrieved items. MOPR is able to balance representation across classes, whereas other baselines miss intersectional groups (highlighted in red).
@@
Figure 8: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 9 additional queries for the CelebA dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 9: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 9 additional queries for the UTKFace dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 10: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 9 additional queries for the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 11: Fraction of Top-$k$ cosine similarity vs. Fraction of Top-$k$ MPR over decision trees for 10 queries for the CelebA dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 12: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over decision trees for 10 queries for the UTKFace dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 13: Fraction of Top-$k$ cosine similarity vs. Fraction of Top-$k$ MPR over decision trees for 10 queries for the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 14: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over 2-layer MLPs (hidden dimension 64) for 10 queries for the CelebA dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 15: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over 2-layer MLPs (hidden dimension 64) for 10 queries for the UTKFace dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 16: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over 2-layer MLPs (hidden dimension 64) for 10 queries for the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 17: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 45 (First 15, see Figures 18 and 19) occupations present in the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 18: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 45 (Second 15, see Figures 17 and 19) occupations present in the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 19: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR over linear regression models for 45 (Third 15, see Figures 17 and 18) occupations present in the Occupations dataset. Values are normalized so Top-$k$ achieves point (1,1) in each case. MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 20: Precision-MPR curve on Occupations dataset over linear regression. Retrieving 50 items for the query “A photo of a chief executive officer." MPR remains competitive in terms of retrieval performance.
@@
Figure 22: Retrievals for MOPR with query “A photo of a teacher"
@@
Figure 25: Retrievals for CLIP-Clip with query “A photo of a teacher"
@@
Figure 28: Retrievals for MOPR  with query “A photo of an architect"
@@
Figure 31: Retrievals for CLIP-Clip with query “A photo of an architect"
@@
Appendix H Experiment Notes
H.1 GPT Generated Queries.
@@ Appendix H Experiment Notes
H.2 Experiment Setting and Details.
These experiments were conducted mainly on CPU after initially using a single A100 GPU to compute CLIP and DebiasCLIP embeddings and run FAISS to compute the top 10k candidates for each. Then, all embeddings are saved so that retrieval across all methods can be run on the CPU. When training linear probes, we conduct a cross-validated hyperparameter search from $0.01$ to $100$ over $\ell_{2}$ regularized logistic regression models from scikit-learn [69] to find the optimal probe.
To run MOPR, we use Gurobi [98] with an Academic License to solve Eqn. (9) each iteration of our cutting plane method and use CVXPY [99, 100] to solve our closed-form, quadratic program.
@@
On average, the runtime for MOPR is approximately 1-5 minutes with 50 queries and 10000 elements in each of our retrieval and curation sets. However, with more features (such as using CLIP embeddings instead of group attributes), this runtime can increase. We note, however, that our most competitive baseline, MMR, is a greedy algorithm that requires 1-2 hours per query to compute. In live image retrieval settings, this runtime is too slow to be of any practical use.